# Run LoRA Run: Faster and Lighter LoRA Implementations

### v1

@@ Abstract
LoRA is a technique that reduces the number of trainable parameters in a neural network by introducing low-rank adapters to linear layers. This technique is used both for fine-tuning (LoRA, QLoRA) and full train (ReLoRA). This paper presents the RunLoRA framework for efficient implementations of LoRA that significantly improves the speed of neural network training and fine-tuning using low-rank adapters. The proposed implementation optimizes the computation of LoRA operations based on dimensions of corresponding linear layer, layer input dimensions and lora rank by choosing best forward and backward computation graph based on FLOPs and time estimations, resulting in faster training without sacrificing accuracy. The experimental results show up to 17% speedup on Llama family of models.
Keywords Low-rank  $\cdot$ LoRA  $\cdot$ Llama
@@ 1 Introduction
LoRA Hu et al. (2022) paper has introduced low-rank adapters to fine-tune large LLMs on downstream tasks. This approach quickly became popular due to reduced cost of the update. Different modifications of LoRA followed, for example, QLoRA Dettmers et al. (2023) utilizes quantization and further reduces fine-tuning costs, and ReLoRA Lialin et al. (2023) which showed that low-rank updates can also be used for full training.
However, all variations of LoRA use the same chain of operations while calculating the output, which often leads to sub-optimal graph of computations. We propose RunLora: a framework which contains different variations of forward and backward pass through an adapter-induced linear layer and chooses the best pair for a given architecture.
We evaluated our framework’s performance on a series of Llama models and achieved up to 17% speedup only due to optimized chain of PyTorch operations. Additionally, we managed to save up to 4Gb of memory due to reduction in number of saved activations.
@@ 2 Problem setting and Methodology
    $\mathrm{d}A=X^{\top}(\mathrm{d}YB^{\top}),\mathrm{d}B=(A^{\top}X^{\top})\mathrm{d}Y,\mathrm{d}X=\mathrm{d}YW^{\top}+(\mathrm{d}YB^{\top})A^{\top}.$
@@
    $\mathrm{d}A=X^{\top}(\mathrm{d}YB^{\top}),\mathrm{d}B=A^{\top}(X^{\top}\mathrm{d}Y),\mathrm{d}X=\mathrm{d}YW^{\top}+(\mathrm{d}YB^{\top})A^{\top}.$
@@
    $\mathrm{d}A=(X^{\top}\mathrm{d}Y)B^{\top},\mathrm{d}B=A^{\top}(X^{\top}\mathrm{d}Y),\mathrm{d}X=\mathrm{d}YW^{\top}+(\mathrm{d}YB^{\top})A^{\top}.$
@@
    $\mathrm{d}A=(X^{\top}\mathrm{d}Y)B^{\top},\mathrm{d}B=A^{\top}(X^{\top}\mathrm{d}Y),\mathrm{d}X=\mathrm{d}Y(W^{\top}+B^{\top}A^{\top}).$
@@
    $\mathrm{d}A=X^{\top}(\mathrm{d}YB^{\top}),\mathrm{d}B=(A^{\top}X^{\top})\mathrm{d}Y,\mathrm{d}X=\mathrm{d}Y(W^{\top}+B^{\top}A^{\top}).$
@@
    $\mathrm{d}A=(X^{\top}\mathrm{d}Y)B^{\top},\mathrm{d}B=(A^{\top}X^{\top})\mathrm{d}Y,\mathrm{d}X=\mathrm{d}YW^{\top}+(\mathrm{d}YB^{\top})A^{\top}.$
@@
    $\mathrm{d}A=(X^{\top}\mathrm{d}Y)B^{\top},\mathrm{d}B=(A^{\top}X^{\top})\mathrm{d}Y,\mathrm{d}X=\mathrm{d}Y(W^{\top}+B^{\top}A^{\top}).$
@@
    $\mathrm{d}A=X^{\top}(\mathrm{d}YB^{\top}),\mathrm{d}B=A^{\top}(X^{\top}\mathrm{d}Y),\mathrm{d}X=\mathrm{d}Y(W^{\top}+B^{\top}A^{\top}).$
However, out of 8 variants we implement only 5 as we assume others to require more or equal number of FLOPs in any setting. For example, backward6 would require more flops than backward5 for any architecture and training configuration (Table 1).
  1. 1.
    backward1, equivalent to the variant 1 above:
    $Z_{1}=\mathrm{d}YB^{\top},Z_{2}=XA,\mathrm{d}A=X^{\top}Z_{1},\mathrm{d}B=Z_{2}^{\top}\mathrm{d}Y,\mathrm{d}X=\mathrm{d}YW^{\top}+Z_{1}A^{\top},$
  2. 2.
    backward2, equivalent to the variant 2 above:
    $Z_{1}=\mathrm{d}YB^{\top},Z_{2}=X^{\top}\mathrm{d}Y,\mathrm{d}A=X^{\top}Z_{1},\mathrm{d}B=A^{\top}Z_{2},\mathrm{d}X=\mathrm{d}YW^{\top}+Z_{1}A^{\top},$
  3. 3.
    backward3, equivalent to the variant 3 above:
    $Z_{1}=\mathrm{d}YB^{\top},Z_{2}=X^{\top}\mathrm{d}Y,\mathrm{d}A=Z_{2}B^{\top},\mathrm{d}B=A^{\top}Z_{2},\mathrm{d}X=\mathrm{d}YW^{\top}+Z_{1}A^{\top},$
  4. 4.
    backward4, equivalent to the variant 4 above:
    $Z_{1}=W+AB,Z_{2}=X^{\top}\mathrm{d}Y,\mathrm{d}A=Z_{2}B^{\top},\mathrm{d}B=A^{\top}Z_{2},\mathrm{d}X=\mathrm{d}YZ_{1}^{\top},$
  5. 5.
    backward5, equivalent to the variant 5 above:
    $Z_{1}=\mathrm{d}YB^{\top},Z_{2}=XA,Z_{3}=W+AB,\mathrm{d}A=X^{\top}Z_{1},\mathrm{d}B=Z_{2}^{\top}\mathrm{d}Y,\mathrm{d}X=\mathrm{d}YZ_{3}^{\top}.$
@@
forward3 $2b\cdot s\cdot(i\cdot o+r\cdot i+o\cdot r)$
forward4 $2(i\cdot o\cdot r+b\cdot s\cdot o\cdot i)$
@@ 3 Numerical experiments
We have conducted a number of experiments on LLama family of models ranging from 60 millions of parameters up to 1.3 billion. We measured mean time of a Forward-Backward pass through the network for different architectures and training settings. As showed in Table 3, we managed to achieve up to 17% speedup compared to default peft implementation.
Additionally, since RunLoRA forward functions do not save intermediate result $XA$ we manage to save up to 4Gb of memory.
All experiments were performed on a single Nvidia A100 GPU 80Gb. Training was conducted in bfloat16 format.
@@
Table 2: Comparison between RunLora (with flops criterion) and default LoRA implementation from peft library. $b$-batch size, $r$-lora rank.

### v2

@@ Abstract
LoRA is a technique that reduces the number of trainable parameters in a neural network by introducing low-rank adapters to linear layers. This technique is used both for fine-tuning and full training of large language models. This paper presents the RunLoRA framework for efficient implementations of LoRA that significantly improves the speed of neural network training and fine-tuning using low-rank adapters. The proposed implementation optimizes the computation of LoRA operations based on dimensions of corresponding linear layer, layer input dimensions and lora rank by choosing best forward and backward computation graph based on FLOPs and time estimations, resulting in faster training without sacrificing accuracy. The experimental results show up to 28% speedup on language modeling networks.
Keywords LoRA  $\cdot$ FLOPs  $\cdot$ Computation graph
@@ 1 Introduction
LoRA Hu et al. (2022) paper has introduced the idea of updating a low-rank correction instead of the full weight matrix. This approach quickly became popular due to reduced cost of the update: number of parameters in the adapter is significantly lower than the original due to a low-rank structure. There emerged a number of papers that proved LoRA’s efficacy not only for fine-tuning on downstream tasks, but also for full training (ReLoRALialin et al. (2023)) or style-transfer (ZipLoRAShah et al. (2023)). Different modifications of LoRA followed as well, incorporating quantization (QLoRADettmers et al. (2023)), weight-sharing (LoTRBershatsky et al. (2024), VeRAKopiczko et al. (2024)), etc..
However, all variations of LoRA use the default chain of operations while calculating the output, which often leads to sub-optimal graph of computations. And none of the papers on low-rank adapter training considers computation costs. We propose RunLora: a framework which contains different variations of forward and backward pass through an adapter-induced linear layer and chooses the best pair for a given architecture. We provide a thorough analysis (both empirically and theoretically) on areas of applicability of each pass in the parameter space.
Our framework is compatible with pytorch and can be used as a simple model wrapper, similar to Lora implementation from PEFT library. We also provide a functionality to work with quantized model weights to fine-tune models in fasion of QLORA Dettmers et al. (2023) paper.
We evaluated our framework’s performance on a series of NLP models, including RoBerta, OPT and Llama, and achieved up to 28% speedup (Figure 1) only due to optimized chain of PyTorch operations. Additionally, we managed to save up to 6Gb of memory due to reduction in number of saved activations (Table 3).
The summary of our contributions is as follows:
  1. 1.
    We implemented a number of alternative forward and backward computation passes through low-rank adapters and investigated areas of applicability of each pass under parameter reduction condition.
  2. 2.
    We developed a framework called RunLoRA: a model wrapper for training with low-rank adapters which uses the best forward-backward passes considering the model architecture and training parameters.
  3. 3.
    We evaluate our framework on a number of common language models: OPT, Llama, RoBerta and demonstrate significant speedups (up to 28%) proving efficiency of RunLora.
Figure 1: Maximum speedups for forward-backward pass through network achieved on different families of models and with different data types. Note that OPT implementation used in these experiments does not support float32 training. b - batch size, r - lora rank, s - sequence length.
@@ 2 Problem setting and Methodology
    $\mathrm{d}A=X^{\top}(\mathrm{d}YB^{\top})$,
    $\mathrm{d}B=(A^{\top}X^{\top})\mathrm{d}Y$,
    $\mathrm{d}X=\mathrm{d}YW^{\top}+(\mathrm{d}YB^{\top})A^{\top}.$
@@
    $\mathrm{d}A=X^{\top}(\mathrm{d}YB^{\top})$,
    $\mathrm{d}B=A^{\top}(X^{\top}\mathrm{d}Y)$,
    $\mathrm{d}X=\mathrm{d}YW^{\top}+(\mathrm{d}YB^{\top})A^{\top}.$
@@
    $\mathrm{d}A=(X^{\top}\mathrm{d}Y)B^{\top}$,
    $\mathrm{d}B=A^{\top}(X^{\top}\mathrm{d}Y)$,
    $\mathrm{d}X=\mathrm{d}YW^{\top}+(\mathrm{d}YB^{\top})A^{\top}.$
@@
    $\mathrm{d}A=(X^{\top}\mathrm{d}Y)B^{\top}$,
    $\mathrm{d}B=A^{\top}(X^{\top}\mathrm{d}Y)$,
    $\mathrm{d}X=\mathrm{d}Y(W^{\top}+B^{\top}A^{\top}).$
@@
    $\mathrm{d}A=X^{\top}(\mathrm{d}YB^{\top})$,
    $\mathrm{d}B=(A^{\top}X^{\top})\mathrm{d}Y$,
    $\mathrm{d}X=\mathrm{d}Y(W^{\top}+B^{\top}A^{\top}).$
@@
    $\mathrm{d}A=(X^{\top}\mathrm{d}Y)B^{\top}$,
    $\mathrm{d}B=(A^{\top}X^{\top})\mathrm{d}Y$,
    $\mathrm{d}X=\mathrm{d}YW^{\top}+(\mathrm{d}YB^{\top})A^{\top}.$
@@
    $\mathrm{d}A=(X^{\top}\mathrm{d}Y)B^{\top}$,
    $\mathrm{d}B=(A^{\top}X^{\top})\mathrm{d}Y$,
    $\mathrm{d}X=\mathrm{d}Y(W^{\top}+B^{\top}A^{\top}).$
@@
    $\mathrm{d}A=X^{\top}(\mathrm{d}YB^{\top})$,
    $\mathrm{d}B=A^{\top}(X^{\top}\mathrm{d}Y)$,
    $\mathrm{d}X=\mathrm{d}Y(W^{\top}+B^{\top}A^{\top}).$
However, out of eight variants we implement only first five (Algorithms 1 - 5) since others require more or equal number of FLOPs in any setting. Specifically, backward6 would require more FLOPs than backward5 for any architecture and training configuration and backward7, backward8 require the same number of FLOPSs as backward3 (Table 1).
Algorithm 1 backward1
$Z_{1}\leftarrow\mathrm{d}YB^{\top}$
$Z_{2}\leftarrow XA$
$\mathrm{d}A\leftarrow X^{\top}Z_{1}$
$\mathrm{d}B\leftarrow Z_{2}^{\top}\mathrm{d}Y$
$\mathrm{d}X\leftarrow\mathrm{d}YW^{\top}+Z_{1}A^{\top}$
Algorithm 2 backward2
$Z_{1}\leftarrow\mathrm{d}YB^{\top}$
$Z_{2}\leftarrow X^{\top}\mathrm{d}Y$
$\mathrm{d}A\leftarrow X^{\top}Z_{1}$
$\mathrm{d}B\leftarrow A^{\top}Z_{2}$
$\mathrm{d}X\leftarrow\mathrm{d}YW^{\top}+Z_{1}A^{\top}$
Algorithm 3 backward3
$Z_{1}\leftarrow\mathrm{d}YB^{\top}$
$Z_{2}\leftarrow X^{\top}\mathrm{d}Y$
$\mathrm{d}A\leftarrow Z_{2}B^{\top}$
$\mathrm{d}B\leftarrow A^{\top}Z_{2}$
$\mathrm{d}X\leftarrow\mathrm{d}YW^{\top}+Z_{1}A^{\top}$
Algorithm 4 backward4
$Z_{1}\leftarrow W+AB$
$Z_{2}\leftarrow X^{\top}\mathrm{d}Y$
$\mathrm{d}A\leftarrow Z_{2}B^{\top}$
$\mathrm{d}B\leftarrow A^{\top}Z_{2}$
$\mathrm{d}X\leftarrow\mathrm{d}YZ_{1}^{\top}$
Algorithm 5 backward5
$Z_{1}\leftarrow\mathrm{d}YB^{\top}$
$Z_{2}\leftarrow XA$
$Z_{3}\leftarrow W+AB$
$\mathrm{d}A\leftarrow X^{\top}Z_{1}$
$\mathrm{d}B\leftarrow Z_{2}^{\top}\mathrm{d}Y$
$\mathrm{d}X\leftarrow\mathrm{d}YZ_{3}^{\top}$
Then we analyse area of applicability for each backward, considering a necessary condition on parameters reduction: number of trainable parameters after LoRA transform should be less than that of the original layer.
$r(i+o)<io$ (5)
where $r$ denotes LoRA rank, $i$ and $o$ denote input and output dimensions respectively.
Figure 2 depicts a case study examples for some batch sizes and sequence lengths. Areas of color depict best choice of forward or backward pass based on minimal number of required FLOPs. Subfigures 2(a) and 2(d) on the left consider a square weight layer where number of input features and output features equal model’s embedding size (i.e., query, key and value layers in transformers). Subfigures 2(b) and 2(c) on the right depict expanding linear layer (typically, 4x expantion is used in MLP blocks of transformers). In all cases, parameter reduction is satisfied only under the dashed line.
In all depicted cases backward2 and backward3 did not appear as the best choice which satisfies condition 5. It can be further proved that neither backward2 nor backward3 will provide the least number of FLOPs under this restriction. It is sufficient to prove that at least one of other backward algorithms is a better option. For both cases, it is convenient to compare against backward5. We will use proof by contradiction:
Suppose FLOPs(backward2) $\leq$ FLOPs(backward5). From 1 it follows that:
$2bs(or+2ir+2io)+2ior$ $\leq 2bs(2or+2ir+oi)+2ior$
$2bsor+4bsir+4bsio+2ior$ $\leq 4bsor+4bsir+2bsoi+2ior$
$2bsio$ $\leq 2bsor$
$i$ $\leq r$
Using 5 and knowing that $i>0,o>0$:
$i\leq r$ $<\frac{io}{i+o}\leq i$
We reached a contradiction. That means FLOPs(backward2) > FLOPs(backward5).
Suppose FLOPs(backward3) $\leq$ FLOPs(backward5). From 1 it follows that:
$2bs(2io+or+ir)+4ior$ $\leq 2bs(2or+2ir+oi)+2ior$
$4bsio+2bsor+2bsir+4iro$ $\leq 4bsor+4bsir+2bsoi+2ior$
$2bsio+2iro$ $\leq 2bsor+2bsir$
$bs(io-or-ir)$ $\leq-iro$
Using 5 and knowing that $i>0,r>0,o>0,b>0,s>0$:
$0<bs$ $\leq\frac{-iro}{io-or-ir}<0$
We reached a contradiction. That means FLOPs(backward3) > FLOPs(backward5).
@@
(a) Input and output dimensions are equal to the model’s embedding size. batch size = 2, sequence length = 100.
(b) Input dimension equals model’s embedding size, output dimension is four times bigger. batch size = 2, sequence length = 100.
(c) Input and output dimensions are equal to the model’s embedding size. batch size = 20, sequence length = 1024.
(d) Input dimension equals model’s embedding size, output dimension is four times bigger. batch size = 1, sequence length = 600.
Figure 2: Areas of best forward/backward pass choice. Region under the dashed line satisfies condition 5.
Areas of applicability can also be researched in batch size/sequence length space. For example, Figure 3 depicts best backward and forward paths for the LlamaMLP linear layer with adapters of rank 128 (this configuration satisfies condition 5).
(a) Best backward path as a function of batch size and sequence length.
(b) Best forward path as a function of batch size and sequence length.
Figure 3: Areas of best forward/backward pass choice for the LlamaMLP linear layer. Input dimention equals 4096, output dimention equals 11008. Rank is 128.
@@ 3 Numerical experiments
To evaluate RunLora performance, we have conducted experiments on several NLP Models with number of parameters ranging from 60 millions up to 1.3 billion: LLama Touvron et al. (2023), OPT Zhang et al. (2022), RoBERTa Liu et al. (2020). We measured mean time of a Forward-Backward pass through the network for different architectures and training settings and compared it to Peft LoRA implementation. Additionally, we performed 100 epochs of adapters training on wikitext-2 Merity et al. (2016) and compared steps-per-second and samples-per-second metrics as well as total training runtime.
Llama
We used Llama model implementation which employs Flash Attention from PyTorch framework. We consider several configurations (from 60 millions of parameters up to 1.3 billions) and initialize layer weights randomly. As showed in Table 3, we managed to achieve up to 16% speedup compared to Peft when running model with float32 data type for weights and operations. When running the same experiment in bfloat16 (Table 3) we manage to achieve up to 17.8% speedup. This slight improvement results from the fact that training in bfloat16 is generally faster than training in full precision, which makes reduction of flops due to RunLoRA to be more influencing on the loop runtime.
When training Llama for 100 epochs on wikitext-2, we manage to achieve 17.56% reduction in total runtime. Accordingly, number of training samples per second and number of train steps per seconds increase in 1.2 times (Table  8).
RoBerta
Another family of models we consider in our experiments consists of roberta-base and roberta-large pretrained models from Huggingface Hub. They contain about 125 millions and 355 millions of parameters respectively. In terms of mean Forward-Backward RunLoRA performs 11.88% faster in float32 and 22.06% faster in bloat16 data type.
As for training RoBerta on wikitext-2, RunLoRA shows up to 20.27% speedup in total runtime and 1.25 times increase in train samples per second and train steps per second (Table 9).
OPT
As with RoBerta, we use pretrained weights and model implementation from Huggingface Hub. With OPT models we also use FlashAttention2 Dao (2024) mechanism which only supports bfloat16 data type.
Table 3 shows maximum of 28.29% speedup of Forward-Backward pass for sequence length of 512 and 26.24% for maximum sequence length of the model. Thereafter, wikitext-2 training experiment (Table 10) depicts maximum 26.65% reduction in total runtime, 1.36 increase in train samples per second and train steps per second.
Additionally, since RunLoRA forward functions do not save intermediate result $XA$, in certain experiments we manage to save up to 6Gb of GPU memory.
Llama2-7b
We also evaluate our framework on a large model with 7 billion parameters. Since training such model on a single gpu proves to be a tedious and often impossible task, we use LitGPT [^1] https://github.com/Lightning-AI/litgpt framework to get advantages of FSDP training. We train Llama2-7b on Alpaca dataset for 100 iteration steps. We use two GPUs with minibatch size of 40. Results are presented in Table 11: RunLoRA achieves 21.47% speedup in mean iteration time.
All experiments were performed on a single Nvidia A100 GPU 80Gb (except for Llama2-7b experiment, which was conducted on two gpus). In all experiments, lora dropout was fixed to 0, other parameters are stated in the referenced tables. For measuring mean Forward-Backward pass we utilized torch.benchmarking [^2] https://pytorch.org/docs/stable/benchmark_utils.html package. RunLoRA adapters were applied to all linear weights in attention and MLP blocks.
Note that we use Adam optimizer Kingma and Ba (2014) and torch.autocast [^3] https://pytorch.org/docs/stable/amp.html when training adapters. Adam requires memory to store optimizer states and autocast converts some activations to fp32 data type. This forces us to use smaller batch sizes or sequence lengths for wikitext-2 training compared to mean Forward-Backward measurements.
Table 2: Comparison between RunLora (with flops criterion) and default LoRA implementation from peft library. $b$-batch size, $r$-lora rank on Llama family of models. Sequence length was set to the maximum sequence length of the model (2048). Experiments were run in float32 data type.
[…]
Table 3: Comparison between RunLora (with flops criterion) and default LoRA implementation from peft library. $b$-batch size, $r$-lora rank on Llama family of models. Sequence length was set to the maximum sequence length of the model (2048). Experiments were run in bfloat16 data type.
[…]
Table 4: Comparison between RunLora (with flops criterion) and default LoRA implementation from peft library. $b$-batch size, $r$-lora rank on OPT family of models. Sequence length was set to the maximum sequence length of the model (2048). Experiments were run in bfloat16 data type.
@@
Table 5: Comparison between RunLora (with flops criterion) and default LoRA implementation from peft library. $b$-batch size, $r$-lora rank on OPT family of models. Sequence length was set to 512. Experiments were run in bfloat16 data type.
@@
Table 6: Comparison between RunLora (with flops criterion) and default LoRA implementation from peft library. $b$-batch size, $r$-lora rank on RoBERTa family of models. Sequence length was set to the maximum sequence length of the model(512). Experiments were run in float32 data type.
[…]
Table 7: Comparison between RunLora (with flops criterion) and default LoRA implementation from peft library. $b$-batch size, $r$-lora rank on RoBERTa family of models. Sequence length was set to the maximum sequence length of the model(512). Experiments were run in bfloat16 data type.
[…]
Table 8: Training Llama Causal LM model for 100 epochs on wikitext-2 dataset. Training was conducted in bfloat16 data type on a single A100 GPU. Texts are grouped to batches with sequence length of 1024 tokens.
Implementation Train Samples per Second Train Steps per Second Train Runtime, Min Speedup, %
llama-130m, b=60, r=128
PEFT 89.46 1.51 51.91 -
RunLoRA 103.9 1.75 44.69 13.91
llama-250m, b=48, r=128
PEFT 47.1 0.99 98.54 -
RunLoRA 56.2 1.19 82.62 16.16
llama-350m, b=40, r=128
PEFT 38.1 0.96 121.98 -
RunLoRA 46.2 1.16 100.56 17.56
llama-1.3b, b=24, r=128
PEFT 11.18 0.47 415.25 -
RunLoRA 12.46 0.52 372.69 10.2
Table 9: Training RoBerta Causal LM model for 100 epochs on wikitext-2 dataset. Training was conducted in bfloat16 data type on a single A100 GPU. Texts are grouped to batches with sequence length of 512 tokens.
Implementation Train Samples per Second Train Steps per Second Train Runtime, Min Speedup, %
roberta-base, b=64, r=128
PEFT 95.84 1.5 85.4 -
RunLoRA 115.04 1.8 71.17 16.7
roberta-large, b=46, r=128
PEFT 42.79 0.93 186.87 -
RunLoRA 53.67 1.18 148.99 20.27
Table 10: Training OPT Causal LM model for 100 epochs on wikitext-2 dataset. Training was conducted in bfloat16 data type on a single A100 GPU. Texts are grouped to batches with sequence length of 1024 tokens.
Implementation Train Samples per Second Train Steps per Second Train Runtime, Min Speedup, %
opt-125m, b=48, r=128
PEFT 75.63 1.6 51.89 -
RunLoRA 96.85 2.1 40.53 21.89
opt-350m, b=32, r=128
PEFT 34.07 1.07 115.19 -
RunLoRA 46.45 1.46 84.49 26.65
opt-1.3b, b=20, r=128
PEFT 15.81 0.79 248.29 -
RunLoRA 20.03 1.0 196.01 21.05
Table 11: Training Llama2-7b model for 100 iterations on Alpaca dataset. Training was conducted in bfloat16 data type on two A100 GPUs. Texts were grouped to batches with sequence length of 512 tokens. Minibatch size = 40. LoRA rank = 128.
Implementation Mean Iteration Time, Sec Speedup, % Train Runtime, Sec Memory used, GB
PEFT 7283.90 - 719.76 24.09
RunLoRA 5720.12 21.47 573.35 23.72